# Beehive: A Flexible Network Stack for Direct-Attached Accelerators

### v3

@@ 6.8. Hardware Resource Utilization
We also evaluate the scalability of the hardware implementation by building a UDP network stack as before and then duplicating the echo application tile. On the Alveo U200, we can place 22 application tiles and 28 tiles total. We are limited by timing rather than resource utilization; the critical path is between NoC routers. Each router is fairly expensive, because the 512-bit width of the bus results in a number of high-fanout wires. This is exacerbated by the fact that the FPGA part in the Alveo U200 is made up of several chiplets, and chiplet crossings add significant delay. Several FPGAs (achronix_noc, ; xilinx_versal, ) now support hardened NoC resources and could improve the quality of results.
7. Related Work
7.1. Packet processing
@@
PANIC (panic_osdi, ) is a framework that supports integration of arbitrary packet processing elements, including general purpose cores. It uses a similar model to Beehive of chaining message-passing elements over a NoC, but it relies on a crossbar, limiting scalability. While the paper does not directly address deadlock, their central scheduler drops packets when it runs out of buffer space, preventing deadlock. This is not suitable for reliable network processing above layer 3. Because it aims to be a NIC framework, PANIC does not need to support higher-level protocols. For example, it does not support multi-packet application requests.
@@
ClickNP (clicknp_sigcomm, ) also supports the integration of arbitrary processing elements. However, it does not use a NoC. Instead components are directly connected via FIFOs, which makes it harder to replicate elements. It also assumes a PCIe connection to a CPU, which it relies on for control plane configuration.
@@
Rosebud (rosebud_asplos, ) is an FPGA framework for middleboxes. It uses an interconnect to connect custom processing elements they call reconfigurable processing units (RPUs) that can include accelerators. Because it targets middleboxes, they do not evaluate a network stack with full reliable transport protocol support. While it does provide support to chain RPUs, they acknowledge it was not designed to do so, and inter-RPU traffic has a fairly significant latency penalty.
@@
A more restrictive approach leverages reconfigurable match-action tables. An action (e.g. strip a header, rewrite a field, drop a packet) is taken based on some header fields in the header of the packet. Typically, there is a pipeline of these processing elements (azure_smartNIC, ; flexnic_asplos, ; rmt_sigcomm, ). However, match-action style processing is not well-suited for highly stateful processing (flowblaze_nsdi, ) typical for application-level offloads. Other models have been proposed for stateful packet processing. Flowblaze uses an FSM-based model (flowblaze_nsdi, ). However, they specifically say that workloads above the transport layer are out of scope. hXDP proposed a processor for eBPF bytecode (hxdp_osdi, ) designed for offloading kernel-level eBPF programs. Because of its sequential execution model, hXDP performs best on small programs and is a poor fit for more complex processing such as Reed-Solomon encoding.
7.2. Transport protocol offloads
@@
Another related vein of work are transport protocol offloads. Most of these are TCP offload engines available as custom chips (chelsio_t6, ) or encrypted IP cores for FPGAs (chevin_toe, ; mle_toe, ; easics_toe, ). They generally do not support the full range of functions found in datacenter network stacks.
@@
Some TCP offload engines could potentially support modification. Limago (limago, ) is an open-source TCP and RoCEv2 offload engine written in Vivado HLS. However, it does not provide any specific APIs or hooks for adding other protocols, so introducing a new network function or new protocol would require fairly extensive modifications to the stack itself. Tonic (tonic, ) is an open-source implementation of the TCP send path and supports customization of the transport protocol, but does not address any lower-level packet processing layers; it also lacks a complete receive path implementation. FlexTOE (flextoe_nsdi, ) is a software implementation of TCP offload engine using the Netronome DPU, a processor designed specifically for network processing that is programmable using C or eBPF. While they do support network functions, their work targets TCP offload for CPUs while our work shows that a direct-attached hardware accelerator does not need a CPU core to support software stack functionality.
@@
Microsoft Catapult’s FPGAs use a custom transport protocol called LTL (catapult_v2, ), which is a reliable transport protocol over UDP. Similar to most TCP engines, it is presented as a fixed IP core with no interface for extension. Catapult also supports a single-layer RMT, used for network virtualization (azure_smartNIC, ). However, it is unknown if these are ever combined and if so, how it would support new protocols or network functions.
8. Conclusion
@@
Modern datacenter networking relies on a variety of network functions and protocols, but current hardware network stacks fall short on these features. As datacenters continue to offload computation to accelerators, it is becoming increasingly important to enable direct-attached accelerators to reduce network overhead. In this paper, we presented the design and implementation of Beehive, a NoC-based network stack for direct-attached accelerators designed to be customizable and to support the variety of protocols and management functions in datacenter networking. We demonstrated that Beehive can combine replicated protocol elements and replicated applications for higher bandwidth, consistent low latency, and minimal overhead. We have open-sourced Beehive for reuse.

### v5

@@ VII-I Scalability
To evaluate hardware resource usage scability, we duplicate echo application tiles connected to a UDP stack. On the Alveo U200, we can place 22 application tiles and 28 tiles total. We are limited by timing rather than resource utilization; the critical path is between NoC routers. Each router is fairly expensive, because the 512-bit width of the bus results in a number of high-fanout wires. This is exacerbated by the fact that the FPGA part in the Alveo U200 is made up of several chiplets, and chiplet crossings add significant delay. Several FPGAs [1, 78] now support hardened NoC resources and could improve the quality of results.
Figure 12: Packet size vs. goodput for a UDP echo application, running on Beehive with multiple network stacks instantiated. 2 network stacks maxes out the load balancer’s throughput.
VIII Related Work
VIII-A Packet processing
@@
PANIC [49] is a smartNIC framework that supports integration of arbitrary packet processing elements, including general purpose cores. Unlike Beehive, PANIC targets packet processing rather than full-stack support for application accelerators. PANIC uses a similar model to Beehive of chaining message-passing elements over a NoC, but it relies on a crossbar, limiting scalability. While PANIC does not directly address deadlocks, its central scheduler drops packets when it runs out of buffer space, preventing deadlock. However, this makes integrating RPC/TCP applications into PANIC is challenging, since it assumes that operations occur on a packet level and the scheduler may drop an acknowledged packet, violating TCP semantics.
@@
ClickNP [47] is an FPGA-accelerated packet processing framework that also supports the integration of arbitrary processing elements. However, it does not use a NoC. Instead, components are directly connected via FIFOs, which makes it harder to replicate elements. Since ClickNP aims to accelerate software network functions, it lacks support for higher-level network protocols and direct-attached accelerators. It further assumes a PCIe connection to a CPU, which it relies on for control-plane configuration.
@@
Rosebud [42] is an FPGA framework for middleboxes. It uses an interconnect to connect custom processing elements they call reconfigurable processing units (RPUs) that can include accelerators. Because it targets middleboxes, they do not evaluate a network stack with full reliable transport protocol support. While it does provide support to chain RPUs, they acknowledge it was not designed to do so, and inter-RPU traffic has a fairly significant latency penalty.
@@
A more restrictive approach leverages reconfigurable match-action tables. An action (e.g. strip a header, rewrite a field, drop a packet) is taken based on some header fields in the header of the packet. Typically, there is a pipeline of these processing elements [30, 40, 11]. However, match-action style processing is not well-suited for highly stateful processing [62] typical of application-level offloads. Other models have been proposed for stateful packet processing. Flowblaze uses an FSM-based model [62]. However, they specifically say that workloads above the transport layer are out of scope. hXDP proposed a processor for eBPF bytecode [12] designed for offloading kernel-level eBPF programs. Because of its sequential execution model, hXDP performs best on small programs and is a poor fit for more complex processing such as Reed-Solomon encoding.
VIII-B Transport protocol offloads
@@
Another related vein of work are transport protocol offloads. Most of these are TCP offload engines available as custom chips [16] or encrypted IP cores for FPGAs [77, 56, 25]. They generally do not support the full range of functions found in datacenter network stacks.
@@
Some TCP offload engines could potentially support modification. Limago [66] is an open-source TCP and RoCEv2 offload engine written in Vivado HLS. However, it does not provide any specific APIs or hooks for adding other protocols, so introducing a new network function or new protocol would require fairly extensive modifications to the stack itself. Tonic [4] is an open-source implementation of the TCP send path and supports customization of the transport protocol, but does not address any lower-level packet processing layers; it also lacks a complete receive path implementation. FlexTOE [69] is a software implementation of TCP offload engine using the Netronome DPU, a processor designed specifically for network processing that is programmable using C or eBPF. While they do support network functions, their work targets TCP offload for CPUs while our work shows that a direct-attached hardware accelerator does not need a CPU core to support software stack functionality.
@@
Microsoft Catapult’s FPGAs use a custom transport protocol called LTL [13], which is a reliable transport protocol over UDP. Similar to most TCP engines, it is presented as a fixed IP core with no interface for extension. Catapult also supports a single-layer RMT, used for network virtualization [30]. However, it is unknown if these are ever combined and if so, how it would support new protocols or network functions.
IX Conclusion
@@
Modern datacenter networking relies on a variety of network functions and protocols, but current hardware network stacks fall short on these features. As datacenters continue to offload computation to accelerators, it is becoming increasingly important to enable direct-attached accelerators to reduce network overhead. In this paper, we presented the design and implementation of Beehive, a NoC-based network stack for direct-attached accelerators that is customizable and supports the variety of protocols and management functions needed for datacenter networking. We demonstrated that Beehive can combine replicated protocol elements and replicated applications for higher bandwidth, provide consistent low latency, with minimal overhead. We have open-sourced Beehive for reuse at https://github.com/beehive-fpga/beehive.
Acknowledgments
We thank the anonymous reviewers for feedback that substantially improved the paper. This work was supported by grants from VMware Research, the National Science Foundation (CNS-2104548, No. 2213387), the University of Washington Center for the Future of Cloud Infrastructure (FOCI), the Intel TSA center, and the NSF Graduate Research Fellowship. Pratyush Patel assisted in an early draft of this paper. We would like to thank Hugo Sadok for his invaluable help with getting Ensō running. We also thank Alexey Lavrov and Jonathan Balkind for feedback on various drafts of this paper.
Artifact Appendix
IX-A Abstract
Beehive is a NoC-based network stack for direct attached accelerators designed to enable flexible construction of complex network functionality in hardware. This appendix outlines the steps to access Beehive. All files necessary to build the various artifacts for the experiments in the paper are available in the Beehive repository.
Full evaluation and reproduction of evaluation results requires testbed access to a 100 Gb switch, at least 4 100 Gb NICs, and an Alveo U200 FPGA. It also requires access to Vivado 2021.2 to build the hardware designs. Simulation requires access to Python and ModelSim 2019.2.
IX-B Artifact check-list (meta-information)
  • Program: Vivado 2021.2, ModelSim 2019.2
  • Compilation: Follow the directions inside beehive/README.md
  • Hardware: Alveo U200/U250
  • How much disk space required (approximately)?: 200 MB
  • Publicly available?: Yes, see below
  • Code licenses (if publicly available)?: BSD 3-Clause
  • Data licenses (if publicly available)?: BSD 3-Clause
  • Archived (provide DOI)?: https://doi.org/10.5281/zenodo.13308868
IX-C Description
IX-C1 How to access
Beehive is accessible at https://github.com/beehive-fpga/beehive
IX-D Installation
Please refer to the instructions inside beehive/README.md
IX-E Evaluation and expected results
Each hardware build needs some sort of input data to run on. You are expected to connect to the hardware via TCP or UDP and send the appropriate packets to drive the device and to fetch evaluation information from the device. If simulating, the simulation driver will drive the design. In this case, you should see outputs relating to that run.
IX-F Methodology
Submission, reviewing and badging methodology: